# A Continuous Energy Ising Machine Leveraging Difference-of-Convex Programming

### v1

@@ 2 Results
We implemented and tested our algorithm across a range of NVIDIA GPUs: 4 GB 128-core Maxwell (Jetson Nano), 4 GB RTX 3050 (Laptop), 24 GB RTX 3090, 32 GB V100, and 80 GB H100. A $60,000$ mAh battery-powered edge computing setup based on the Jetson Nano module is shown in Supplementary Figure 12. We tested the effectiveness of our algorithm on the MAX-CUT problem, where the task is to partition the vertices of a graph into two disjoint subsets that maximize the number of edges between them [39]. This NP-hard problem is commonly used as a benchmark for evaluating Ising solvers. We also implemented and compared our method with several state-of-the-art algorithms such as Simulated Annealing (SA) [17, 31], ballistic Bifurcation Machine (bSB) [27], Simulated Coherent Ising Machine (SimCIM) [31], and Spring Ising Algorithm (SIA) [38]. For completeness, detailed descriptions of these algorithms, along with the procedure for constructing ultra-large-scale Ising models, are provided in the Supplementary Notes 7 to 9.
@@
We benchmark various Ising solvers on the Sherrington-Kirkpatrick (SK) model [49], a fully connected Ising model where the symmetric coupling coefficients $J_{ij}=J_{ji}\,(i\neq j)$ are sampled from the standard normal distribution $\mathcal{N}(0,1)$. Each solver is tasked with minimizing the Ising energy of a system with $10^{3}$ spins. Their performance is evaluated based on the time required to reach the energy level obtained by the Goemans-Williamson Semidefinite Program (GW-SDP). As the solutions in the SK model typically converge quickly, we restrict the runtime of each solver to under one second. All experiments are repeated for $100$ random initializations. As shown in Figure 4a, ADOCH reaches the GW-SDP energy threshold in about $15$ ms. Moreover, the histograms in Figures 4b and 4c show that ADOCH consistently finds the lowest energy states with the highest frequency. Overall, our algorithms demonstrate the most robust and optimal performance at steady state compared to the tested methods.
Figure 4: Evolution of Ising energy with runtime for the $10^{3}$-spin SK model. (a) Comparison of our solvers (DOCH and ADOCH) with state-of-the-art Ising solvers: Simulated Annealing (SA) [17, 31], ballistic Bifurcation Machine (bSB) [27], Simulated Coherent Ising Machine (SimCIM) [31], and Spring Ising Algorithm (SIA) [38]. Solid curves show the best energy values achieved across 100 independent trials. The dashed black line indicates the lowest energy obtained using the GW-SDP [26, 31] with 100 random rounding projections (Supplementary Notes 6 and 7 for the parameter settings). (b), (c) Histograms of Ising energies from 100 runs of each algorithm after 1000 iterations. The results were obtained on a laptop equipped with an NVIDIA RTX 3050 GPU.
@@
We next evaluate our algorithm on the standard $K_{2000}$ benchmark, a fully connected graph with $2000$ nodes and nearly $2$ million edges that is commonly used for MAX-CUT problems. The MAX-CUT problem can be formulated as an Ising model, where the coupling matrix is given by $\mathbf{J}=-(1/2)\mathbf{W}$, with $\mathbf{W}$ denoting the weighted adjacency matrix of the graph (Supplementary Note 5). For our experiments,the weights $\{W_{ij}\}$ are sampled independently from $\{-1,1\}$ with equal probability. Since the ground state is not available, we benchmark solvers based on the lowest Ising energy $\mathcal{E}$ achieved within a fixed runtime of 10 s (excluding data loading) and evaluate their consistency across $100$ random initializations. We also compare the time each solver takes to reach the energy level obtained using GW-SDP. As shown in Figure 5a, the DOCH and ADOCH solvers outperform the best-performing spring Ising algorithm (SIA), reaching the GW-SDP benchmark in approximately $50$ and $70$ ms. While SIA slightly outpaces DOCH in reaching the GW-SDP energy level, DOCH ultimately achieves the lowest Ising energy within the full $10$-second window. Figures 5b and 5c further demonstrate that both DOCH and ADOCH consistently yield superior solutions across multiple runs with different initializations.
@@
We again exploit the equivalence between the Ising model and the MAX-CUT problem to solve MAX-CUT on the $G_{10}$ graph, an 800-node, $94.01\%$ sparse graph with $\pm 1$ edge weights from the G-set dataset. As shown in Figure 7a, our algorithms consistently outperform other heuristic Ising solvers. Although each solver is run for $1000$ iterations, DOCH and ADOCH take only about $100$ iterations to reach the best solution. In particular, ADOCH reaches GW-SDP-level performance within just $3$ iterations, followed by DOCH and SIA.
@@
We also evaluate the solvers on random graphs with integer-valued edge weights from the MAX-CUT benchmark suite in the Biq Mac Library [50]. Each solver is run for $1000$ iterations and over $100$ different initializations. The best cut values are shown as a bar graph in Figure 6b. We used the cut value obtained using $10^{5}$ iterations of SA as the largest cut value. For each run, we measure the time taken to reach $99\%$ of the largest cut value and average this across multiple runs to compute the average time-to-solution (Avg TTS). As shown in Figure 6a, the DOCH, ADOCH solvers consistently achieve the lowest Avg TTS across all graph instances, demonstrating superior convergence speed. Figure 6b further shows that DOCH and ADOCH attain the best or comparable cut values compared to other methods. All experiments on these small-scale Ising problems were conducted using a Jetson Nano module, except for the $K_{2000}$ and $1000$ spin SK model, which were run in a laptop equipped with an NVIDIA RTX 3050 GPU.
Figure 5: Ising energy versus runtime on the $K_{2000}$ benchmark. (a) Comparison of our solvers (DOCH and ADOCH) with state-of-the-art Ising solvers. Solid curves show the mean Ising energy across 100 independent trials. The dashed black line indicates the lowest energy obtained using the GW-SDP [26, 31] with 100 random rounding attempts. (b), (c) Histograms of Ising energies from 100 runs of each algorithm after 1000 iterations. The results were obtained on a laptop equipped with an NVIDIA RTX 3050 GPU.
Figure 6: MAX-CUT results on Biq Mac graphs. (a) Table summarizing the graph characteristics and the average time-to-solution (Avg TTS) for each solver to reach 99% of the maximum cut value. (Best Avg TTS values are highlighted in red, and second-best in blue). (b) Bar chart comparing the best cut values achieved within 1000 iterations by SA [17, 31], bSB [27], SimCIM [31], SIA [38], and our Ising solvers DOCH and ADOCH. The results were obtained on NVIDIA Jetson Nano.
@@
For Ising models of this scale, GW-SDP becomes computationally impractical to run. Moreover, for models with $10^{4}$ spins, we require more powerful GPUs such as RTX 3090 and V100. In Figure 7b, we illustrate the performance of various Ising solvers on the $10^{4}$ spin SK model. Each algorithm was run for 20 s with the same initialization until saturation. The plot in Figure 7b shows the time evolution of the Ising energy for each solver. We observe that ADOCH outperforms the other solvers, achieving an energy below $5\times 10^{5}$ in under 0.5 s. DOCH also performs competitively, reaching a comparable energy level around 5 s and converging close to the value achieved by ADOCH. Performance comparisons for larger Ising models with $10^{5}$ and $10^{6}$ spins are provided in (Supplementary Figures 4-9).
We also compare our Ising solver with the recently introduced free energy machine (FEM) [51], which is based on the principle of free-energy minimization. FEM shares some similarities with our approach, notably in its use of continuous relaxation and the incorporation of an external entropy term in the energy function. However, FEM inherits limitations common to SA methods, including sensitivity to hyperparameters, reliance on a carefully designed cooling schedule, and challenges in parameter tuning. Moreover, FEM employs optimizers like RMSprop and ADAM [52] which do not offer convergence guarantees. Additionally, gradient computation becomes a significant bottleneck in FEM, complicating its implementation for large-scale problems ($n>10^{4}$). As shown in (Supplementary Figures 1 and 2), our DOCH solvers consistently outperform FEM within 1 second for both small ($10^{3}$ spin) and medium ($10^{4}$ spin) SK models. It is also important to acknowledge that FEM is a more generalized formalism applicable to both the MAX-CUT and multi-spin Ising problems.
@@
We benchmark the Ising solvers on ultra-large-scale Ising model problems involving $10^{7}-10^{8}$ spins, covering both sparse and dense connectivity regimes. Figure 7c shows the performance on a fully connected Ising model with $10^{7}$ spins with nearly $50$ trillion couplings. The coupling coefficients are generated using the pseudo-random function: $J_{ij}=\sin(i\,j+\mathrm{seed})$ with $\mathrm{seed}=100$. The experiments were performed using four NVIDIA H100 GPUs. The runtime versus Ising energy plot in Figure 7c clearly demonstrates the superior performance of our methods. In particular, the ADOCH achieves the lowest energy levels with the fastest convergence, followed closely by DOCH. These results underscore the scalability and efficiency of our approach in tackling both sparse and fully connected Ising models at unprecedented scales. Although the execution of Ising solvers on fully connected graphs demands much more computational resources than sparsely connected graphs.
We further demonstrate the performance of the Ising solvers on a $10^{8}$ spin and $0.00001\%$ connected Ising model, having approximately 0.5 billion nonzero coupling coefficients. The nonzero $J_{ij}$ are sampled uniformly from the set of 9-bit signed integers. The matrix computations were performed using two H100 GPUs with parallelized matrix-vector multiplication (Supplementary Note 10). Each algorithm was run for $10^{3}$ s ($\approx 17$ minutes), during which the Ising energy reached saturation. The resulting runtime versus Ising energy plots are shown in Figure 7d. Both DOCH and ADOCH demonstrate superior performance, achieving the lowest energy values among all solvers. Additional results on similarly ultra-large-scale Ising models are provided in (Supplementary Figures 10 and 11). To our knowledge, no prior work has reported Ising solvers at this scale.
Figure 7: Benchmarking solver performance at varying Ising model sizes. (a) MAX-CUT optimization on an 800-node $G_{10}$ graph. The dashed black line marks the lowest energy found using the GW-SDP (with $100$ random rounding trials). All algorithms were executed on an NVIDIA Jetson Nano, with cut values plotted against iterations. (b) Benchmarking on a $10^{4}$-spin Sherrington–Kirkpatrick (SK) model, executed on NVIDIA Jetson Nano. Ising energy is shown as a function of time (log scale) in seconds. (c) Fully connected $10^{7}$-spin Ising model with couplings $J_{ij}=\sin(i\,j+\mathrm{seed})$. Experiments were conducted using four NVIDIA H100 GPUs. The energy trajectories over time are shown. (d) Sparse $10^{8}$-spin Ising model with $0.00001\%$ connectivity and 99-bit signed integer couplings ($J_{ij}\in\{-2^{9}+1,\dots,2^{9}-1\}$), executed on two NVIDIA H100 GPUs. Energy is plotted against computation time (in minutes).
@@ 3 Discussion
We introduced our Ising solver by relaxing the binary spins into continuous variables and formulating the resulting optimization as a difference-of-convex program. This approach was motivated by the success of first-order optimization methods for large-scale continuous optimization, particularly in modern deep learning [52]. Central to our method is a tunable attractor function, which is coupled with the Ising energy to guide solutions toward binary spin assignments. Exploiting the structure of this attractor, we designed two simple yet effective iterative algorithms (DOCH and ADOCH), requiring just a single matrix-vector multiplication per iteration. We further established that DOCH is guaranteed to converge to a critical point of the Hamiltonian. In contrast to conventional Ising solvers, our approach avoids reliance on cooling schedules or extensive hyperparameter tuning, making it especially well-suited for ultra-large-scale Ising problems.
To demonstrate scalability, we solved a fully connected Ising model with $10^{7}$ spins and $\sim 50$ trillion couplings (Figure 7c), significantly surpassing the size of previously reported models, which maxed out at $10^{6}$ spins and $\sim 5$ billion couplings [27]. On smaller but dense graphs such as $K_{2000}$, DOCH performs best over short time horizons (1-10 s), while ADOCH exhibits faster convergence for large to ultra-large-scale models. For the benchmarks including $K_{2000}$, $10^{3}$-spin SK model and Biq Mac graphs, we ran our solvers with 100 random initializations. The histograms in Figures 4b, 4c, 5b, and 5c demonstrate that our methods are robust to the initialization. Our solvers rapidly achieve GW-SDP-level energy values and consistently produce the highest cut-values on the $G_{10}$ graph and Biq Mac instances. Overall, our solvers exhibit faster convergence to lower-energy configurations across the board, including ultra-large-scale models; see Figures 7c and 7d, and also (Supplementary Figures 3-11).
@@
In summary, our solvers offer a compelling alternative to traditional simulation-based, cooling-dependent approaches. Their simple update rules allow for efficient implementation on GPU clusters, enabling scalability to ultra-large problem instances. Furthermore, the inherently parallel structure of the algorithms makes them well-suited for deployment on low-level hardware such as FPGAs, offering significant gains in both computational speed and energy efficiency.
@@ 4 Methods
First-order methods have become the cornerstone of large-scale optimization, particularly in deep learning and related fields [52, 53, 54]. Their success is driven by a combination of low computational cost per iteration and the ability to scale efficiently to problems with extremely large parameter spaces. By relying solely on gradient information, these methods bypass the need to compute or invert Hessians, making them highly effective for nonconvex problems involving millions or even billions of variables. Additionally, their algorithmic simplicity allows for straightforward parallelization and efficient deployment on hardware accelerators, which has been critical for scaling optimization in modern, data-intensive applications.
Our Ising solver builds on the observation that the Hamiltonian in (2) can be expressed as a difference of convex functions. A particularly effective technique for optimizing such objectives is the Difference-of-Convex Algorithm (DCA)[40], which is rooted in the classical principle of bound optimization[46]. In this approach, the original nonconvex objective is iteratively approximated by a sequence of convex surrogate problems that upper-bound the original function. These surrogates are much easier to solve and are particularly amenable to first-order methods, making DCA highly suitable for large-scale problems like the Ising model.
As we explain next, applying DCA to our problem (9) results in a particularly simple iterative scheme. Starting from an initial point $\boldsymbol{x}^{(k)}$, we linearize the concave component g-g using its first-order approximation around $\boldsymbol{x}^{(k)}$. This yields a convex surrogate, which is minimized to obtain the next iterate $\boldsymbol{x}^{(k+1)}$. In particular, as $g$ is convex, its linear approximation at $\boldsymbol{x}^{(k)}$ gives us a global lower bound [54]. Specifically, for all $\boldsymbol{x}\in\mathbb{R}^{n}$,
@@
We then refine our estimate by minimizing this upper bound. Specifically, we define the surrogate function:
$F(\boldsymbol{x})=f(\boldsymbol{x})-g(\boldsymbol{x}^{(k)})-\nabla\!g(\boldsymbol{x}^{(k)})^{\top}\!(\boldsymbol{x}-\boldsymbol{x}^{(k)}),$
and set $\boldsymbol{x}^{(k+1)}$ as its minimizer. Since $F$ is convex and differentiable, using first-order optimality, we get
$\nabla F(\boldsymbol{x}^{(k+1)})=\nabla\!f(\boldsymbol{x}^{(k+1)})-\nabla\!g(\boldsymbol{x}^{(k)})=\mathbf{0}.$ (12)
@@
That is, $\mathcal{H}$ is guaranteed to decrease or remain the same at each iteration. Since $\mathcal{H}$ is both continuous and coercive, the descent property (14) ensures convergence to a limiting value. Furthermore, because $\mathcal{H}$ is a polynomial function, we can establish that the sequence of iterates produced by DOCH is stable, in the sense that $\{\boldsymbol{x}^{(k)}\}$ converges to a limit point $\boldsymbol{x}^{*}\in\mathbb{R}^{n}$. We additionally prove that this limit $\boldsymbol{x}^{*}$ is (unconditionally) a critical point of the Hamiltonian $\mathcal{H}$, and under mild assumptions, is a strict local minimizer (Supplementary Note 3).
Algorithm 1 DOCH
@@
A natural extension of our algorithm is to incorporate acceleration into DCA [41]. Accelerated first-order methods improve the convergence speed of traditional first-order algorithms by incorporating a momentum mechanism. A prominent example is Nesterov’s acceleration [47, 52], which achieves the optimal convergence rate for convex problems, outperforming standard gradient descent. Accelerated methods preserve the simplicity and low per-iteration cost of standard first-order methods, while achieving faster convergence. This makes them especially effective for large-scale optimization problems. In our case, we apply Nesterov-style acceleration to develop the Accelerated DOCH (ADOCH) algorithm, with the main steps outlined in Algorithm 4.
Algorithm 2 Accelerated DOCH (ADOCH)
@@
The key distinction from DOCH is the incorporation of momentum:
@@
where we extrapolate $\boldsymbol{x}^{(k)}$ using the previous iterate $\boldsymbol{x}^{(k-1)}$. The parameter $t_{k}$ is set following Nesterov’s optimal first-order scheme [47]. To decide whether to accept the extrapolated point $\boldsymbol{y}^{(k)}$, we evaluate the condition:
$\mathcal{H}(\boldsymbol{y}^{(k)})\leqslant\max\big{\{}\mathcal{H}(\boldsymbol{x}^{(k-q)}),\dots,\mathcal{H}(\boldsymbol{x}^{(k)})\big{\}}$ (15)
where $q$ controls the look-back window [41]. This Barzilai–Borwein type scheme helps the solver to escape multiple local minima of $\mathcal{H}$ [55]. Theoretically, a large value of $q$ ensures better acceleration and quicker convergence [56]. If condition (15) is met, we apply the update $\boldsymbol{x}^{(k+1)}=\mathcal{T}(\boldsymbol{y}^{(k)})$; otherwise, we revert back to the non-extrapolated point. This alternate update at $\boldsymbol{y}^{(k)}$ instead of $\boldsymbol{x}^{(k)}$ is the principal difference with DOCH.

### v2

@@ 2 Results
We have implemented and tested our algorithms across a range of NVIDIA GPUs, including a 4 GB 128-core Maxwell processor (Jetson Nano), 4 GB RTX 3050 (laptop), 24 GB RTX 3090, 32 GB V100, and 80 GB H100. A battery-powered 60,000 mAh edge-computing setup based on the Jetson Nano module is shown in Supplementary Figure 12. We have tested the effectiveness of our algorithms on the MAX-CUT problem, where the task is to partition the vertices of a graph into two disjoint subsets that maximize the number of edges between them [goemans1995]. This NP-hard problem is commonly used as a benchmark for evaluating Ising solvers. We have also implemented and compared our method with several state-of-the-art algorithms such as Simulated Annealing (SA) [isakov2015, inagaki2016], ballistic Bifurcation Machine (bSB) [goto2021], Simulated Coherent Ising Machine (SimCIM) [goto2021, inagaki2016], and Spring Ising Algorithm (SIA) [jiang2024]. To avoid being trapped in local minima, we have used $100$ different initializations while running each solver on small to medium-scale Ising models. For completeness, detailed descriptions of these algorithms, along with the procedure for constructing ultra-large-scale Ising models, are provided in Supplementary Notes 7 to 9.
@@
We benchmark various Ising solvers on the Sherrington-Kirkpatrick (SK) model [sherrington1975], a fully connected Ising model where the symmetric coupling coefficients $J_{ij}=J_{ji}\,(i\neq j)$ are sampled from the standard normal distribution $\mathcal{N}(0,1)$. Each solver is tasked with minimizing the Ising energy of a system with $10^{3}$ spins. Their performance is evaluated based on the time required to reach the energy level obtained by the GW-SDP (using $100$ random rounding projections). As the solutions in the SK model typically converge quickly, we restrict the runtime of each solver to under $0.1$ s (excluding the time needed to load data and compute the Ising energy). All experiments are repeated for $100$ random initializations. As shown in Figure 4a, DOCH and ADOCH reach the GW-SDP energy threshold in about $2.5$ and $3.5$ ms respectively. Moreover, the histograms in Figures 4b and 4c show that ADOCH consistently finds the lowest energy states with the highest frequency. Overall, our algorithms demonstrate the most robust and optimal performance at steady state compared to the other solvers.
@@
Next, we evaluate our algorithm on the standard $K_{2000}$ benchmark, a fully connected graph with $2000$ nodes and nearly $2$ million edges that is commonly used for MAX-CUT problems. The MAX-CUT problem can be formulated as an Ising model, where the coupling matrix is given by $\mathbf{J}=-(1/2)\mathbf{W}$, with $\mathbf{W}$ denoting the weighted adjacency matrix of the graph (Supplementary Note 5). For our experiments, the weights $\{W_{ij}\}$ are sampled independently from $\{-1,1\}$ with equal probability. Since the ground state is not available, we benchmark solvers based on the lowest Ising energy $\mathcal{E}$ achieved within a fixed runtime of $1$ s (excluding data loading and energy calculation) and evaluate their consistency across $100$ random initializations. We also compare the time each solver takes to reach the energy level obtained by GW-SDP. As shown in Figure 5a, the DOCH and ADOCH solvers outperform the best-performing spring Ising algorithm (SIA), reaching the GW-SDP benchmark in approximately $1.9$ and $2.6$ ms respectively. DOCH on average performs the best within the one-second window. Figures 5b and 5c further demonstrate that both DOCH and ADOCH consistently yield superior solutions across multiple runs with different initializations.
@@
We solve MAX-CUT on the (random) $G_{10}$ graph, an $800$-node, $94.01\%$ sparse graph with $\pm 1$ edge weights from the G-set dataset. We compute the mean cut values over $100$ random initializations as shown in Figure 7a. Each solver is run for $1000$ iterations. DOCH and ADOCH take only about $100$ iterations to reach the best solution. In particular, ADOCH attains GW-SDP level performance after just $3$ iterations on average, closely followed by DOCH and SIA.
Results on additional G-set graphs are provided in Supplementary Figure 12. There, our solvers maintain superior performance with consistently smaller error bars even after $100$ iterations, especially on random $(G_{6},G_{7},G_{8})$ and toroidal $(G_{11},G_{12},G_{13})$ graphs. For longer runs ($\geqslant 100$ iterations), bSB and SimCIM gradually improve and beat our solvers for planner graphs $G_{18},G_{19},G_{20},G_{21}$. However, such long runs are computationally prohibitive for ultra-large models. For example, on $10^{7}$-spin fully connected graphs, other solvers will take about $1000$ hours to run for $100$ iterations, including hyperparameter tuning (see Ultra-large models section).
@@
We also evaluate the solvers on random graphs with integer-valued edge weights from the MAX-CUT benchmark suite in the Biq Mac Library [bigq_mac]. Each solver is run for $1000$ iterations and over $100$ different initializations. The best cut values are shown as a bar graph in Figure 6b. We used the cut value obtained after $10^{5}$ iterations of SA as the largest cut value. For each run, we measure the time taken to reach $99\%$ of that largest cut value and average this across multiple runs to compute the average time-to-solution (average TTS). As shown in Figure 6a, the DOCH, ADOCH solvers consistently achieve the lowest Avg TTS across all graph instances, demonstrating superior convergence speed. Figure 6b further shows that DOCH and ADOCH attain the best or comparable cut values with respect to the other solvers.
The experiments on these small-scale Ising problems are conducted using a Jetson Nano module, except for the $K_{2000}$, the 1000-spin SK model, and the G-set graphs shown in Supplementary Figure 13. The computations are performed on a NVIDIA RTX 3050 GPU laptop with 4 GB RAM.
@@
For Ising models of this scale, GW-SDP becomes computationally impractical to run. Moreover, for models with $10^{4}$ spins, we require more powerful GPUs such as RTX 3090 and V100. In Figure 7b, we illustrate the performance of various Ising solvers on the $10^{4}$ spin SK model. Each algorithm is run for $1$ s with the same initialization until saturation. The plot in Figure 7b shows the time evolution of the Ising energy for each solver. We observe that ADOCH outperforms other solvers, and DOCH achieves an energy below $5\times 10^{5}$ in under $0.1$ s. Performance comparisons for larger Ising models with $10^{5}$ and $10^{6}$ spins are provided in (Supplementary Figures 4-7).
We also compare our Ising solver with the recently introduced free energy machine (FEM) [fem2025], which is based on the principle of free-energy minimization. FEM shares some similarities with our approach, notably in its use of continuous relaxation and the incorporation of an external entropy term in the energy function. However, FEM inherits limitations common to SA methods, including sensitivity to hyperparameters, reliance on a carefully designed cooling schedule, and challenges in parameter tuning. Moreover, FEM employs optimizers like RMSprop and ADAM [2015-kingma] which do not offer convergence guarantees. Additionally, gradient computation becomes a significant bottleneck in FEM, complicating its implementation for large-scale problems ($n>10^{4}$). As shown in (Supplementary Figures 1 and 2), our DOCH solvers consistently outperform FEM within $0.1$ and $1$ s for small ($10^{3}$ spin) and medium ($10^{4}$ spin) SK models respectively. It is also important to acknowledge that FEM is a more generalized formalism applicable to both the MAX-CUT and multi-spin Ising problems.
@@
We benchmark the Ising solvers on ultra-large-scale Ising model problems involving $10^{7}-10^{8}$ spins, covering both sparse and dense connectivity regimes. Figure 7c shows the performance on a fully connected Ising model with $10^{7}$ spins with nearly $50$ trillion couplings. The coupling coefficients are generated using the pseudo-random function: $J_{ij}=\sin(ij+\mathrm{seed})$ with $\mathrm{seed}=100$. The experiments are performed using four NVIDIA H100 GPUs. The runtime versus Ising energy plot in Figure 7c clearly demonstrates the superior performance of our methods. In particular, the ADOCH achieves the lowest energy level, followed closely by DOCH. These results underscore the scalability and efficiency of our solvers in tackling both sparse and fully connected Ising models at unprecedented scales. However the execution of Ising solvers on fully connected graphs demands much more computational resources than sparsely connected graphs.
With a fixed runtime budget, state-of-the-art solvers must complete a full run for every hyperparameter choice to achieve the best Ising energy, while our solvers only need to inspect the first few iterations, allowing much faster parameter tuning. For instance, in the $10^{7}$-spin ultra-large Ising model from Figure 7c, tuning 10 hyperparameter values requires about $100$ hours for other solvers but only $20$ hours for ours. Subsequently, running each solver for $11$ hours showed that DOCH surpassed all others within $8$ hours, while ADOCH achieved its best results within $11$ hours.
We further demonstrate the performance of the Ising solvers on a $10^{8}$-spin and $0.00001\%$ connected Ising model, having approximately 0.5 billion nonzero coupling coefficients. The nonzero $J_{ij}$ are sampled uniformly from the set of 9-bit signed integers. The matrix computations are performed using two H100 GPUs with parallelized matrix-vector multiplication (Supplementary Note 10). Each algorithm is run for $10^{3}$ s ($\approx 17$ minutes), during which the Ising energy reached a saturation point. The resulting runtime versus Ising energy plots are shown in Figure 7d. Both DOCH and ADOCH demonstrate superior performance, achieving the lowest energy values among all solvers. Additional results on similarly ultra-large-scale Ising models are provided in (Supplementary Figures 10 and 11). To our knowledge, no prior work has reported Ising solvers at this scale.
@@ 3 Discussion
We have introduced our Ising solver by relaxing the binary spins into continuous variables and formulating the resulting optimization as a difference-of-convex program. This approach is motivated by the success of first-order optimization methods for large-scale continuous optimization, particularly in modern deep learning [2015-kingma]. Central to our method is a tunable attractor function, which is coupled with the Ising energy to guide solutions toward binary spin assignments. By exploiting the structure of this attractor, we design two simple yet effective iterative algorithms (DOCH and ADOCH), which require only a single matrix-vector multiplication per iteration. We further establish that DOCH is guaranteed to converge to a critical point of the Hamiltonian. In contrast to conventional Ising solvers, our approach avoids reliance on cooling schedules or extensive hyperparameter tuning, making it especially well-suited for ultra-large-scale Ising problems. Our methods require fewer iterations to reach high-quality Ising energies, which leads to a substantially lower overall runtime, even though the per-iteration cost is comparable to competing solvers (see Remarks S1, S2 in the Supplementary).
The key observations from our benchmarking are discussed below:
  • We demonstrate scalability by solving a fully connected Ising model with $10^{7}$ spins and $\sim 50$ trillion couplings (Figure 7c), far exceeding previously reported models capped at $10^{6}$ spins and $\sim 5$ billion couplings [goto2021].
  • On smaller dense graphs such as $K_{2000}$, DOCH performs best over short time horizons, while ADOCH converges fastest for large and ultra-large models.
  • For benchmarking on small-scale Ising models, we used $100$ random initializations; the histograms in Figures 4b, 4c, 5b, and 5c demonstrate strong robustness to initialization.
  • Our solvers rapidly achieve GW-SDP-level energies and consistently perform better on the $G_{10}$ graph and Biq Mac instances.
  • Across all scales, including the ultra-large models in Figures 7c, 7d and Supplementary Figures 1 to $1$, our methods exhibit faster convergence to lower-energy configurations than competing solvers.
@@
In summary, our solvers offer a compelling alternative to traditional simulation-based, cooling-dependent approaches. Their simple update rules enable efficient implementation on GPU clusters, allowing for scalability to ultra-large problem instances. Furthermore, the inherently parallel structure of the algorithms makes them well-suited for deployment on low-level hardware such as FPGAs, offering significant gains in both computational speed and energy efficiency.
@@ 4 Methods
First-order methods have become the cornerstone of large-scale optimization, particularly in deep learning and related fields [2015-kingma, lecun2015deep, wir_beck_first-order_2017]. Their success is driven by a combination of low computational cost per iteration and the ability to scale efficiently to problems with extremely large parameter spaces. By relying solely on gradient information, these methods bypass the need to compute or invert Hessians, making them highly effective for nonconvex problems involving millions or even billions of variables. Additionally, their algorithmic simplicity allows for straightforward parallelization and efficient deployment on hardware accelerators, which has been critical for scaling optimization in modern, data-intensive applications.
Our Ising solver builds on the observation that the Hamiltonian in (2) can be expressed as a difference of convex functions. A particularly effective technique for optimizing such functions is the Difference-of-Convex Algorithm (DCA) [abbaszadehpeivasti2023], which is rooted in the classical principle of bound optimization [dempster1977maximum]. In this approach, the original nonconvex objective is iteratively approximated by a sequence of convex surrogate problems that upper-bound the original function. By designing the surrogate problems to be easier to solve (for example, using efficient first-order methods), we can scale DCA to large-scale problems.
As explained next, applying DCA to our problem (9) results in a simple iterative scheme. Starting from an initial point $\boldsymbol{x}^{(k)}$, we linearize $g$ using its first-order approximation around $\boldsymbol{x}^{(k)}$. This yields a convex surrogate, which is minimized to obtain the next iterate $\boldsymbol{x}^{(k+1)}$. In particular, as $g$ is convex, its linear approximation at $\boldsymbol{x}^{(k)}$ gives us a global lower bound [wir_beck_first-order_2017]. Specifically, for all $\boldsymbol{x}\in\mathbb{R}^{n}$, we have
@@
We refine our current estimate $\boldsymbol{x}^{(k)}$ by minimizing this upper bound. Specifically, we set the update $\boldsymbol{x}^{(k+1)}$ as the minimizer of the surrogate function:
$F(\boldsymbol{x})=f(\boldsymbol{x})-g(\boldsymbol{x}^{(k)})-\nabla\!g(\boldsymbol{x}^{(k)})^{\top}\!(\boldsymbol{x}-\boldsymbol{x}^{(k)}).$
As $F$ is convex and differentiable, by first-order optimality,
$\mathbf{0}=\nabla F(\boldsymbol{x}^{(k+1)})=\nabla\!f(\boldsymbol{x}^{(k+1)})-\nabla\!g(\boldsymbol{x}^{(k)}).$ (12)
@@
That is, $\mathcal{H}$ is guaranteed to decrease or stay constant after each iteration. Since $\mathcal{H}$ is continuous and coercive, the descent property (14) ensures that the sequence $\{\mathcal{H}(\boldsymbol{x}^{(k)}\}$ converges to a limiting value. Moreover, because $\mathcal{H}$ is a polynomial, we can establish that the sequence of iterates $\{\boldsymbol{x}^{(k)}\}$ converges to a limit point $\boldsymbol{x}^{*}\in\mathbb{R}^{n}$. We (unconditionally) prove that $\boldsymbol{x}^{*}$ is a critical point of $\mathcal{H}$, and under mild additional assumptions, that it is a strict local minimizer of $\mathcal{H}$ (Supplementary Note 3).
Algorithm 1: DOCH
@@
A natural extension of our algorithm is to incorporate acceleration into DCA [phan2018]. Accelerated first-order methods improve the convergence speed of traditional first-order algorithms. A prominent example is Nesterov’s acceleration [Nesterov2018, 2015-kingma], which achieves the optimal convergence rate for convex problems, outperforming standard gradient descent. Accelerated methods preserve the simplicity and low per-iteration cost of standard first-order methods, while achieving faster convergence. This makes them especially effective for large-scale optimization problems. In our case, we apply Nesterov-style acceleration to develop the Accelerated DOCH (ADOCH) algorithm, with the main steps outlined in Algorithm 2.
Algorithm 2: Accelerated DOCH (ADOCH)
@@
The main difference with DOCH is the use of momentum. We extrapolate $\boldsymbol{x}^{(k)}$ using the previous update $\boldsymbol{x}^{(k-1)}$:
@@
where the sequence $\{t_{k}\}$ is set using Nesterov’s optimal scheme [Nesterov2018]. To decide whether to accept the extrapolated point $\boldsymbol{y}^{(k)}$, we check the condition
$\mathcal{H}(\boldsymbol{y}^{(k)})\leqslant\max\big\{\mathcal{H}(\boldsymbol{x}^{(k-q)}),\dots,\mathcal{H}(\boldsymbol{x}^{(k)})\big\},$ (15)
where $q$ controls the look-back window [phan2018]. This Barzilai–Borwein type criterion helps the method escape shallow local minima of $\mathcal{H}$ [grippo2002]. Theoretically, a large value of $q$ ensures better acceleration and quicker convergence [wright2009]. If (15) holds, we update $\boldsymbol{x}^{(k+1)}=\mathcal{T}(\boldsymbol{y}^{(k)})$; otherwise, we fall back to the non-extrapolated update. Taking the update at $\boldsymbol{y}^{(k)}$ rather than $\boldsymbol{x}^{(k)}$ is the key distinction from DOCH.